Solve the given initial-value problem. x = (-: 1. )x. xco) - (-:) X(t) =
Answer #1

Solne Given, initial value problem. ac - 6 2 xo) 1 -1 8 12 let A = 6 .X X -1 9 12 حقا . 22 and x= 24 Sot X2 bes Ax. In ordeor (6-1) (12-1) +9 = 0 = 0, or or 72-18dtd tg 42 -18d tal = 0. 08 27 ( A - 9.1 O. de g, g. Repeated eigenvalues, let u = 내 U2Eigenvectors. ur [] 3 1. Let U= V be the another eigenvetor Corresponding to eigenvalue dag where وفا to Then we have А 19 -Cenit ut Coretzt. Date eigenvector v - Bt - 1. ti فا =#3 + - 1 ve -1 + 3 1. 4 Therefore, Generals solution of given System.Date Page Tutti ing values of C and (₂ in (5) we get alt) 8e94 Ano [ 1. 3 1 +25eat {E}] tt 3 1. which is the required Solutio